# Vulnerability-Aware Spatio-Temporal Learning for Generalizable and Interpretable Deepfake Video Detection

### v2

@@ Abstract
Detecting deepfake videos is highly challenging due to the complex intertwined spatial and temporal artifacts in forged sequences. Most recent approaches rely on binary classifiers trained on both real and fake data. However, such methods may struggle to focus on important artifacts, which can hinder their generalization capability. Additionally, these models often lack interpretability, making it difficult to understand how predictions are made. To address these issues, we propose FakeSTormer, offering two key contributions. First, we introduce a multi-task learning framework with additional spatial and temporal branches that enable the model to focus on subtle spatio-temporal artifacts. These branches also provide interpretability by highlighting video regions that may contain artifacts. Second, we propose a video-level data synthesis algorithm that generates pseudo-fake videos with subtle artifacts, providing the model with high-quality samples and ground truth data for our spatial and temporal branches. Extensive experiments on several challenging benchmarks demonstrate the competitiveness of our approach compared to recent state-of-the-art methods. The code is available at https://github.com/10Ring/FakeSTormer.
@@ 1 Introduction
With the advances in generative modeling, deepfake videos have become alarmingly realistic. Despite their interest in several applications such as entertainment and education, this type of technology also raises societal concerns [52, 44, 5]. There is therefore an urgency for developing effective deepfake detection methods.
@@
For that reason, researchers have started investigating video deepfake detection methods that are capable of modeling spatio-temporal artifacts [22, 20, 65, 62, 24, 60, 55]. Those methods mainly rely on a deep neural network formed by a single binary classification branch that is trained using a fixed dataset with real and fake data (see Figure 1-a). As a result, they suffer from two main limitations, namely: (1) The lack of generalizability - As highlighted in [39, 42, 32, 47], models trained with a standalone binary classifier tend to overfit the type of deepfakes they are trained on, resulting in poor generalization to unseen manipulations. To address this issue, multi-task learning and data synthesis strategies have demonstrated their relevance in image-based deepfake detection [32, 47, 6, 63, 42]. Nevertheless, such mechanisms have been disregarded in the field of video-level deepfake detection. (2) The lack of interpretability - As discussed in [18], binary outputs are opaque and difficult to interpret, despite the importance of explaining the decision of deepfake detectors. It has been shown in image-based deepfake detection that adding auxiliary non-binary branches can help in providing more explainable outputs [6, 42, 38, 32]. However, as mentioned earlier, multi-task learning strategies remain underexplored in the field of video-based deepfake detection.
A data synthesis strategy has been explored for video deepfake detection. STC [33] uses a Spatio-Temporal Generator (STG) to create pseudo-fake data from real data, improving model generalizability by blending two sequences, one of which has its frames shuffled in the temporal domain. However, this frame shuffling can introduce noticeable temporal artifacts (illustrations are shown in the supplementary materials). As demonstrated in image deepfake detection [47], pseudo-fake data with subtler artifacts improves generalization. To address this, we propose a method that generates subtler temporal artifacts. Additionally, our multi-task approach provides interpretability for the model’s output.
@@
In this paper, to build a model that is both generalizable and interpretable, we propose utilizing a multi-branch network that trains on synthesized data and integrates learning objectives specifically targeting both spatial and subtle temporal artifacts. As shown in Figure 1-b, a novel multi-task learning framework, termed FakeSTormer is introduced and formed by two auxiliary parallel branches in addition to the standard classification head, namely: (1) a regression temporal branch that aims at predicting the vulnerability-prone temporal locations. It has been demonstrated that regressing spatial vulnerabilities in specific points [42] or patches [41] can help improve the generalizability of a deepfake detector model. We refer to the definitions given in [41, 42] which describe: “vulnerable patches/points as the patches/points that are the most likely to embed blending artifacts”. To generalize this concept to the temporal domain, we propose locating temporal high changes in spatial vulnerable patches (see Figure 1-c). (2) a spatial branch to ensure a balance between the spatial and the temporal domains. In fact, detecting spatial artifacts in addition to temporal ones is crucial [55, 62]. For that purpose, we propose predicting frame-wise spatial vulnerabilities.
@@
To create ground truths for these two branches, we introduce a video-level data synthesis algorithm, called “Self-Blended Videos (SBVs)”, inspired by “Self-Blended Images (SBIs) [47]. Our experiments demonstrate that simply training a baseline classification model on SBVs enables achieving on par performance w.r.t. state-of-the-art (SOTA), highlighting the effectiveness of SBVs. Finally, for enhancing spatial and temporal modeling, we revisit the TimeSformer [2] architecture that we use as our backbone. In particular, we leverage TimeSformer’s decomposed temporal and spatial attention on embedded patches, appending classification tokens for each frame and for each patch across frames, rather than a single token for the entire video. These classification tokens are then used within the spatial and classification heads, while the embedded patches are used within the temporal head. Extensive experiments on several well-known deepfake detection benchmarks show that our method is competitive against the existing SOTA approaches in terms of generalization capabilities, while providing more interpretable outputs through the proposed auxiliary branches.
@@
  • A novel multi-task learning framework for video-based deepfake detection, enhancing both generalization and interpretability aspects.
  • Two auxiliary branches that capture both temporal and spatial vulnerabilities, that are fined-grained.
  • A video-level data synthesis technique called SBVs that generates high-quality deepfakes and is supported by a vulnerability-driven cutout augmentation strategy to avoid overfitting specific artifact-prone regions.
@@
Figure 2: Overview of the proposed framework. Our multi-task learning framework, FakeSTormer consists of three branches, i.e., the temporal branch $h$, the spatial branch $g$, and the standard classification branch $f$. Those branches are specially designed to facilitate disentanglement learning of spatial-temporal features. The ground-truth data to train the framework are generated based on our proposed video-level data synthesis algorithm coupled with a vulnerability-driven Cutout strategy.
2 Related Work: Video-based Deepfake Detection
Using a naive spatio-temporal binary classifier for deepfake video detection can lead to overfitting on obvious artifacts, reducing generalization [65, 55]. To address this, FTCN [65] uses a fully temporal convolution network, reducing the spatial kernel size to minimize focus on spatial artifacts. LipForensics [24] focuses on the mouth region, and spatiotemporal dropout [60] removes parts of input frames across both spatial and temporal domains. AltFreezing [55] convolution layers into spatial and temporal ones, but it can fail to model long-term dependencies. ISTVT [62] uses a Vision Transformer with self-attention (SA) for longer-range correlations, splitting SA into spatial and temporal. Additionally, [9] decomposes features into spatial and temporal components. Differently, TALL [56] employs an image-level deepfake detector by converting video frames into a thumbnail layout.
While promising, the aforementioned methods rely on training binary classifiers with both real and fake data, which can lead to overfitting on fake data artifacts. Training only on real data, as suggested in image-based deepfake detection [32, 6, 42], can improve generalization. Using only real data, the recent STC [33] method generates pseudo-fake samples using time-shuffling, but this creates obvious temporal artifacts. Inspired by image-based methods [47], we propose a pseudo-fake generation technique that creates subtler artifacts, improving generalization. Additionally, our approach incorporates a multi-task setup, which enhances both model performance and interpretability.
@@ 3 Methodology
Let $\mathcal{V}\triangleq\cup_{i=1}^{N}\{(\mathbf{X}_{i},y_{i})\}$ be a training dataset formed by $N$ videos, where $\mathbf{X}_{i}$ denotes the $i^{\text{th}}$ video sample and $y_{i}$ its associated label indicating whether the clip is real ($y_{i}=0$) or fake ($y_{i}=1$). Traditional methods [55, 22, 56, 24, 65, 62, 60, 20] aim to learn jointly a feature extractor $\Phi:\mathcal{V}\mapsto\mathcal{F}$ and a binary classifier $f:\mathcal{F}\rightarrow\{0,1\}$ by minimizing the standard binary cross-entropy (BCE) loss $\mathcal{L}_{BCE}(f(\Phi(\mathbf{X}_{i})),y_{i})$ using the entire training set $\mathcal{V}$, with $\mathcal{F}$ being the learned feature space. As highlighted in Section 1, such a strategy might lead to poor generalization capabilities to unseen generation methods [32, 42, 39, 47] while providing only binary outputs that are not interpretable.
To tackle these issues, inspired by the literature on image-level deepfake detection [32, 42, 6, 63, 15], we introduce a novel multi-task learning framework called FakeSTormer that only relies on the real data subset denoted as $\mathcal{V}^{r}\subset\mathcal{V}$. Specifically, in addition to the binary classifier $f$, our framework includes two additional branches $h:\mathcal{F}\rightarrow\mathcal{H}$ and $g:\mathcal{F}\rightarrow\mathcal{G}$ that aim at triggering the learning of localized temporal and spatial artifact-prone features, respectively, through relevant auxiliary tasks. Note that $\mathcal{G}$ and $\mathcal{H}$ denote respectively the output spaces of $h$ and $g$. The proposed branches are depicted further in Section 3.2. To provide ground truth to those branches and at the same time avoid overfitting to specific manipulations, we apply to each video belonging to $\mathcal{V}^{r}$ a data synthesis method described in Section 3.1, resulting in a pseudo-fake subset denoted as $\mathcal{V}^{\tilde{r}}$. Hence, our framework is trained using $\mathcal{\tilde{V}}\triangleq\{\mathcal{V}^{r}\cup\mathcal{V}^{\tilde{r}}\}$.
Figure 3: a) Overview of generating a self-blended video. It contains two main components including a landmark interpolation module and the consistent utilization of synthesized parameters (CSP). b) Examples of pseudo-fake videos. (with(w/) and without(w/o)) vulnerability-driven Cutout and their corresponding soft labels. We apply the Cutout data augmentation at the same spatial locations throughout video frames.
@@ 3.1 Video-Level Data Synthesis and Augmentation
Self-Blended Videos.
Blending-based data synthesis methods have demonstrated great performance in image-based deepfake detection [47, 32, 42, 6, 7, 63, 41]. In fact, as the blending step is common to different manipulation types, they contribute to the improvement of the generalization aspect in deepfake detection [42, 32]. Nevertheless, such an approach has been overlooked in the context of video-based deepfake detection. Hence, we propose to extend blending-based data synthesis to the video level. In particular, we revisit SBI [47] given its ability to produce high-quality pseudo-fake images that have contributed to improving the generalization capabilities of deepfake detectors. The proposed data synthesis approach termed Self-Blended Videos (SBVs) is constituted of two main components, i.e., a Landmark Interpolation module followed by a Consistent Synthesized Parameters (CSP) module for preserving the temporal coherence of synthesized videos, which is essential for producing high-quality synthesized videos. Specifically, given a real video $\mathbf{X}_{i}^{r}\in\mathcal{V}^{r}$ formed by $T$ consecutive frames, we start by extracting a set of 2D landmarks $\mathbf{L}_{i}(t)=\{\mathbf{l}_{ij}(t)\}_{1\leq j\leq n}$ at each instant $t$ from $\mathbf{X}_{i}^{r}(t)$, where $n$ refers to the number of landmarks and $\{\mathbf{l}_{ij}(t)\}\in\mathbb{R}^{2}$. Then, we apply SBI to the $1^{\text{st}}$ video frame denoted as $\mathbf{X}^{r}_{i}(t_{0})$ to obtain a pseudo-fake image, i.e., $\mathbf{X}_{i}^{\tilde{r}}(t_{0})$ and a blending mask $\mathbf{M}_{i}(t_{0})$. All related blending parameters $\mathbf{\theta}^{(sbi)}$ (e.g., ConvexHull type, Mask deformation kernels, blending ratio, etc.) are then conserved for synthesizing the remaining video frames. However, using those parameters solely cannot guarantee the temporal consistency of pseudo-fake videos since the geometry of landmarks can significantly vary over time. To mitigate the issue, we propose to re-interpolate each landmark $\mathbf{l}_{i}(t)$ based on $\mathbf{l}_{i}(t-1)$ for $t>t_{0}$ as follows,
@@
As a result, we obtain the pseudo-fake video $\mathbf{X}_{i}^{\tilde{r}}\in\mathcal{V}^{\tilde{r}}$ and its blending mask sequence $\mathbf{M}_{i}\in\mathbb{R}^{T\times H\times W}$, with $H$ and $W$ being the image height and width, respectively. An illustration of SBVs is given in Figure 3. Additional samples as well as the detailed algorithm are provided in supplementary materials. It is important to note that despite being simple, SBVs is generic and applicable to any existing video-level deepfake detection approach.
Figure 4: Extraction of temporal vulnerabilities. We compute derivatives of the spatial vulnerabilities over time.
Vulnerability-Driven Cutout Augmentation. Previous works [66, 53] have demonstrated that deep learning methods are often impacted by overfitting. Deepfake detectors might even be more sensitive to this phenomenon as deepfakes are typically characterized by localized artifacts [42]. One solution to regularize training is data augmentation. As such, we propose in addition to SBVs, a novel Cutout data augmentation driven by vulnerable patches, i.e., image patches that are prone to blending artifacts [41]. We posit that by masking the most vulnerable regions, overfitting risks will be reduced, as the model will be pushed to learn from other areas. This masking strategy has already been explored in other computer vision fields such as Classification [12, 8], Object Detection [21] demonstrating great potential.
@@
After that, we define a threshold $\tau_{cutout}$ that is randomly selected from the range $(0.5,1.0]$. We use the latter to define the set of patches to be masked $\mathcal{P}=\left\{(l,m)\mid\bar{\mathbf{B}}_{l,m}(t_{0})>\tau_{cutout}\right\}$ within the first frame. The set $\mathcal{P}$ is then used to mask out patches at those locations not only in the first frame but also in the entire video to enforce temporal consistency that is crucial for generating high-quality pseudo-fakes. After masking those patches over time, we finally obtain the masked blending boundary denoted as $\tilde{\mathbf{B}}$. This results in masking the most vulnerable regions, i.e., the regions that are the most likely to include blending artifacts. Figure 3-b shows some examples of the proposed cutout augmentation.

### v3

# Vulnerability-Aware Spatio-Temporal Learning for Generalizable Deepfake Video Detection

@@ Abstract
Detecting deepfake videos is highly challenging given the complexity of characterizing spatio-temporal artifacts. Most existing methods rely on binary classifiers trained using real and fake image sequences, therefore hindering their generalization capabilities to unseen generation methods. Moreover, with the constant progress in generative Artificial Intelligence (AI), deepfake artifacts are becoming imperceptible at both the spatial and the temporal levels, making them extremely difficult to capture. To address these issues, we propose a fine-grained deepfake video detection approach called FakeSTormer that enforces the modeling of subtle spatio-temporal inconsistencies while avoiding overfitting. Specifically, we introduce a multi-task learning framework that incorporates two auxiliary branches for explicitly attending artifact-prone spatial and temporal regions. Additionally, we propose a video-level data synthesis strategy that generates pseudo-fake videos with subtle spatio-temporal artifacts, providing high-quality samples and hand-free annotations for our additional branches. Extensive experiments on several challenging benchmarks demonstrate the superiority of our approach compared to recent state-of-the-art methods. The code is available at https://github.com/10Ring/FakeSTormer.
@@ 1 Introduction
With the advances in generative modeling [22, 43], deepfake videos have become alarmingly realistic. Despite their interest in several applications, such as entertainment and education, this type of technology also raises societal concerns [53, 45, 5]. There is therefore an urgency for developing effective deepfake detection methods.
@@
For that reason, researchers have explored video-based deepfake detection methods capable of modeling spatio-temporal artifacts [23, 20, 71, 67, 25, 64, 56]. Those methods mainly rely on a deep neural network formed by a single binary classification branch that is trained using a fixed dataset with real and fake data (see Figure 1-a). As a result, they suffer from two main limitations, namely: (1) The lack of generalizability - As highlighted in [39, 42, 33, 48], models trained with a standalone binary classifier tend to overfit the type of deepfakes they are trained on, resulting in poor generalization to unseen manipulations; (2) The lack of robustness to high-quality (HQ) deepfake videos - The quality of deepfake videos is improving continuously, resulting in subtle spatio-temporal artifacts. As such, vanilla single-branch architectures trained solely with binary supervision fail to fully capture them, necessitating the design of appropriate attention mechanisms.
To address the generalization issue, video-level data synthesis approaches [34, 24, 62] have been introduced to encourage models to learn more generic representations. However, these methods usually simulate exaggerated temporal variations that are inherently different from artifacts in hyper-realistic deepfake videos. On the other hand, to model localized spatio-temporal inconsistencies, some recent methods [67, 63] have used dedicated architectures integrating an implicit attention mechanism. Nevertheless, these models still rely solely on a binary classifier with no guarantee of extracting artifact-prone fine-grained traces.
Interestingly, in image-based deepfake detection, it has been recently demonstrated that the use of a tailored multi-task learning framework for explicitly attending artifact-prone small regions coupled with a subtle data synthesis strategy can be a way to enhance generalization and, at the same time, robustness to high-quality deepfakes  [42, 41]. Nevertheless, such an approach has been disregarded in the field of video-level deepfake detection, as its extension to the video level is not straightforward. In particular, it would necessitate the characterization of subtle temporal artifacts that are inherently different from spatial ones, within both the multi-task learning framework and the data synthesis.
@@
In this paper, we redefine deepfake video detection as a fine-grained detection task by proposing a multi-branch network that leverages synthesized data and incorporates specialized learning objectives specifically targeting both subtle spatial and temporal artifacts. As shown in Figure 1-b, a novel multi-task learning framework, termed FakeSTormer is introduced. It is formed by two auxiliary parallel branches in addition to the standard classification head, namely: (1) a regression temporal branch incorporating an explicit attention that aims at locating the vulnerability-prone temporal locations. It has been shown that regressing spatial vulnerabilities in specific points [42] or patches [41] can help improve the generalizability of a deepfake detector model. We refer to the definitions given in [41, 42] which describe: “vulnerable patches/points as the patches/points that are the most likely to embed blending artifacts". To generalize this concept to the temporal domain, we propose locating temporal high changes in spatial vulnerable patches (see Figure 1-c). (2) a spatial branch to ensure a balance between the spatial and the temporal domains. In fact, detecting spatial artifacts in addition to temporal ones is crucial [56, 67]. For that purpose, we propose predicting frame-wise spatial vulnerabilities.
@@
To create hand-free ground truths for the proposed branches, we introduce a HQ video-level data synthesis algorithm, called “Self-Blended Video (SBV)”, inspired by “Self-Blended Image (SBI) [48]”, enforcing temporal coherence using two proposed modules on top of SBI (detailed in Section. 3.1). Our experiments demonstrate that simply training a baseline classification model on SBV enables achieving on par performance w.r.t. state-of-the-art (SOTA), highlighting the effectiveness of SBV. Finally, for enhancing spatial and temporal modeling, we revisit the TimeSformer [2] architecture that we use as our backbone. In particular, we leverage TimeSformer’s decomposed temporal and spatial attention on embedded patches, appending classification tokens for each frame and for each patch across frames, rather than a single token for the entire video. These classification tokens are then used within the spatial and classification heads, while the embedded patches are used within the temporal head. Extensive experiments on several well-known deepfake detection benchmarks show that our method outperforms the existing SOTA approaches.
@@
  • A novel multi-task learning framework using only real data for fine-grained video-based deepfake detection.
  • Two auxiliary branches that capture both temporal and spatial vulnerabilities, that are fined-grained by definition.
  • A video-level data synthesis technique called SBV that generates high-quality pseudo-fakes and is supported by a vulnerability-driven cutout augmentation strategy to avoid overfitting specific artifact-prone regions.
@@
2 Related Work
Video-based Deepfake Detection. As highlighted in [71, 56], using a naive spatio-temporal binary classification model for video-level deepfake detection can lead the model to overfit obvious artifacts, resulting in poor generalization to unseen manipulations. To address this, FTCN [71] proposes a fully temporal convolution network by reducing the spatial kernel size to one, hence decreasing the likelihood of focusing only on spatial artifacts. LipForensics [25] considers solely the mouth region, while spatio-temporal dropout [64] randomly removes parts of the input frames in both spatial and temporal domains. AltFreezing [56] separates convolution layers into spatial and temporal ones, failing to model long-term dependencies. Instead of using convolution layers, ISTVT [67] utilizes a video-based Vision Transformer [16] with self-attention to extract longer-range correlations. Meanwhile, [9] decomposes features into spatial and temporal components. TALL [57] employs an image-level deepfake detector by converting video frames into a thumbnail layout. Despite being promising, most of the aforementioned methods solely rely on a single binary classifier that implicitly guides the feature extraction. As highlighted in the literature on image-based deepfake detection [33, 6, 42, 15], this approach might lead to overfitting specific artifacts present in training datasets. Moreover, the absence of an explicit attention mechanism to spatio-temporal artifact-prone regions can lead to poor robustness to high-quality artifacts.
Data Synthesis. A highly effective approach for enhancing the generalizability of deepfake detectors is training models with synthesized data. While frame-level solutions have been extensively studied [48, 33, 68, 6, 15], video-level augmentations remains relatively underexplored. In recent works, STC [34] generates pseudo-fake samples via time-shuffling, VB [62] perturbs landmarks per frame without imposing temporal coherence, while ST-SBV [24] injects temporal artifacts through random face scaling and blurring over time. However, these methods often introduce exaggerated temporal distortions that differ from HQ deepfakes typically exhibiting finer temporal inconsistencies.
Figure 2: I) Overview of the proposed framework: Our multi-task learning framework, FakeSTormer, consists of three branches, i.e., the temporal branch ($h$), the spatial branch ($g$), and the standard classification branch ($f$). Those branches are specially designed to facilitate the disentanglement learning of spatial-temporal features. The hand-free ground-truth data to train the framework are generated based on our proposed video-level data synthesis algorithm coupled with a vulnerability-driven Cutout strategy. II) Overview of generating a self-blended video: It contains two main components, including a landmark interpolation module (LI) and the consistent utilization of synthesized parameters (CSP). III) Examples of pseudo-fake videos: with(w/) and without(w/o) vulnerability-driven Cutout and their corresponding soft labels. We apply the Cutout data augmentation at the same spatial locations throughout video frames. IV) Extraction of temporal vulnerabilities: We compute derivatives of the spatial vulnerabilities over time.
@@ 3 Methodology
Let $\mathcal{V}\triangleq\cup_{i=1}^{N}\{(\mathbf{X}_{i},y_{i})\}$ be a training dataset formed by $N$ videos, where $\mathbf{X}_{i}$ denotes the $i^{\text{th}}$ video sample and $y_{i}$ its associated label indicating whether the clip is real ($y_{i}=0$) or fake ($y_{i}=1$). Traditional methods [56, 23, 57, 25, 71, 67, 64, 20] aim to learn jointly a feature extractor $\Phi:\mathcal{V}\mapsto\mathcal{F}$ and a binary classifier $f:\mathcal{F}\rightarrow\{0,1\}$ by minimizing the standard binary cross-entropy (BCE) loss $\mathcal{L}_{BCE}(f(\Phi(\mathbf{X}_{i})),y_{i})$ using the entire training set $\mathcal{V}$, with $\mathcal{F}$ being the learned feature space. As previously discussed in [33, 42, 39, 48] and also highlighted in Section 1, such a strategy might lead to poor generalization capabilities to unseen generation methods while providing only binary outputs that are not interpretable.
To tackle these issues, inspired by the literature on image-level deepfake detection [33, 42, 6, 68, 15], we introduce a novel multi-task learning framework called FakeSTormer that only relies on the real data subset denoted as $\mathcal{V}^{r}\subset\mathcal{V}$. Specifically, in addition to the binary classifier $f$, our framework includes two additional branches $h:\mathcal{F}\rightarrow\mathcal{H}$ and $g:\mathcal{F}\rightarrow\mathcal{G}$ that aim at triggering the learning of localized temporal and spatial artifact-prone features, respectively, through relevant auxiliary tasks. Note that $\mathcal{H}$ and $\mathcal{G}$ denote respectively the output spaces of $h$ and $g$. The proposed branches are depicted further in Section 3.2. To provide ground truth to those branches and at the same time avoid overfitting to specific manipulations, we apply to each video belonging to $\mathcal{V}^{r}$ a data synthesis method described in Section 3.1, resulting in a pseudo-fake subset denoted as $\mathcal{V}^{\tilde{r}}$. Hence, our framework is trained using $\mathcal{\tilde{V}}\triangleq\{\mathcal{V}^{r}\cup\mathcal{V}^{\tilde{r}}\}$.
@@ 3.1 Video-Level Data Synthesis and Augmentation
Self-Blended Video.
Blending-based data synthesis methods have demonstrated great performance in image-based deepfake detection [48, 33, 42, 6, 7, 68, 41]. In fact, as the blending step is common to different manipulation types, they contribute to the improvement of the generalization aspect in deepfake detection [42, 33]. Nevertheless, such an approach has been overlooked in the context of video-based deepfake detection. Hence, we propose to extend blending-based data synthesis to the video level. In particular, we revisit Self-Blended Image (SBI) [48], given its ability to produce high-quality pseudo-fake images. The proposed data synthesis approach, termed Self-Blended Video (SBV), is constituted of two main components building on top of SBI, i.e., a Consistent Synthesized Parameters (CSP) module followed by a Landmark Interpolation module (LI) for preserving the temporal coherence of synthesized videos, which is essential for producing high-quality synthesized videos.
Specifically, given a real video $\mathbf{X}_{i}^{r}\in\mathcal{V}^{r}$ formed by $T$ consecutive frames, we start by extracting a set of 2D landmarks $\mathbf{L}_{i}(t)=\{\mathbf{l}_{ij}(t)\}_{1\leq j\leq n}$ at each instant $t$ from $\mathbf{X}_{i}^{r}(t)$, where $n$ refers to the number of landmarks and $\{\mathbf{l}_{ij}(t)\}\in\mathbb{R}^{2}$. Then, we apply SBI to the $1^{\text{st}}$ video frame denoted as $\mathbf{X}^{r}_{i}(t_{0})$ to obtain a pseudo-fake image, i.e., $\mathbf{X}_{i}^{\tilde{r}}(t_{0})$ and a blending mask $\mathbf{M}_{i}(t_{0})$. All related blending parameters $\mathbf{\theta}^{(sbi)}$ (e.g., ConvexHull type, Mask deformation kernels, blending ratio, etc.) are then conserved for synthesizing the remaining video frames. However, using those parameters solely cannot guarantee the temporal consistency of pseudo-fake videos since the geometry of landmarks can significantly vary over time. To mitigate the issue, we propose to re-interpolate each landmark $\mathbf{l}_{i}(t)$ based on $\mathbf{l}_{i}(t-1)$ for $t>t_{0}$ as follows,
@@
As a result, we obtain the pseudo-fake video $\mathbf{X}_{i}^{\tilde{r}}\in\mathcal{V}^{\tilde{r}}$ and its blending mask sequence $\mathbf{M}_{i}\in\mathbb{R}^{T\times H\times W}$, with $H$ and $W$ being the image height and width, respectively. An illustration of SBV is given in Figure 2-II. Additional samples, as well as the detailed algorithm, are provided in supplementary materials. It is important to note that, despite being simple, SBV is generic and applicable to any existing video-level deepfake detection approach.
Vulnerability-Driven Cutout Augmentation. Previous works [72, 54] have demonstrated that deep learning methods are often impacted by overfitting. Deepfake detectors might even be more sensitive to this phenomenon as deepfakes are typically characterized by localized artifacts [42]. One solution to regularize training is data augmentation. As such, we propose, in addition to SBV, a novel Cutout data augmentation driven by vulnerable patches, i.e., image patches that are prone to blending artifacts [41]. We posit that by masking the most vulnerable regions, overfitting risks will be reduced, as the model will be pushed to learn from other areas. This masking strategy has already been explored in other computer vision fields such as Classification [12, 8], Object Detection [21] demonstrating great potential.
@@
After that, we define a threshold $\tau_{cutout}$ that is randomly selected from the range $(0.5,1.0]$. We use the latter to define the set of patches to be masked $\mathcal{P}=\left\{(l,m)\mid\bar{\mathbf{B}}_{l,m}(t_{0})>\tau_{cutout}\right\}$ within the first frame. The set $\mathcal{P}$ is then used to mask out patches at those locations not only in the first frame but also in the entire video to enforce temporal consistency that is crucial for generating high-quality pseudo-fakes. After masking those patches over time, we finally obtain the masked blending boundary denoted as $\tilde{\mathbf{B}}$. This results in masking the most vulnerable regions, i.e., the regions that are the most likely to include blending artifacts. Figure 2-III shows some examples of the proposed cutout augmentation.